OpenStudy (anonymous):

add. show steps. -7/45 + 19/45=

OpenStudy (anonymous):

-7/45 + 19/45 = (-7 + 19) / 45 = 12/45

OpenStudy (anonymous):

thats the answer I got

OpenStudy (anonymous):

Doesnt that need to be simplified?

OpenStudy (anonymous):

could simply to 4/15

OpenStudy (anonymous):

and that answer is incorrect

OpenStudy (anonymous):

okay

OpenStudy (anonymous):

thanks

OpenStudy (anonymous):

you're welcome